OpenStudy (anonymous):

she rides a bike 10 inches her brother rides kiddy sized bike wheels are the radius of 4 what were the difference of the two bikes in one complete rotation? give answer in of (pi)

OpenStudy (gorv):

one complete revolution means diatance=2*pi*r

OpenStudy (gorv):

r= radius

OpenStudy (gorv):

she covered=2*pi*10=20pi her brother covered=2*pi*4=8pi

OpenStudy (gorv):

take their diffference ??